OpenStudy (winsteria):

MEDAL AND FAN Solve. 8 – 3x – x = 12 x=?

OpenStudy (winsteria):

@DanJS Can you check my answer?

OpenStudy (anonymous):

1x

OpenStudy (mathmate):

What is your answer?

OpenStudy (winsteria):

-1?

OpenStudy (anonymous):

3x + 2x - 8 = 12 5x - 8 = 12 5x = 20 x = 4

OpenStudy (mathmate):

That is correct. 8 – 3x – x = 12 Group like terms: 8-12=3x+x -4=4x -1=x

OpenStudy (anonymous):

This is tricky

OpenStudy (anonymous):

-1=x

OpenStudy (winsteria):

Yay thank you

OpenStudy (mathmate):

You're welcome! :) BTW, @siearra your piggy-backed post for help might very likely ignored. You'd be better off starting a new thread.

OpenStudy (anonymous):

I need help in something

OpenStudy (anonymous):

in what @1rarity1

OpenStudy (anonymous):

math